Walt Hubble is contemplating selling rental property that originally cost $200,000. He believes that it has appreciated in value at an annual rate of 6% over its 4-year holding period. He will have to pay a commission equal to 5% of the sale price to sell the property. Currently, the property has a book value of $137,000. The mortgage balance outstanding at the time of sale currently is $155,000. Walt will have to pay a 15% tax on any capital gains and a 25% tax on recaptured depreciation.

a. Calculate the tax payable on the proposed sale.

b. Calculate the after-tax net proceeds associated with the proposed sale, CFR.
Answer #1

a.

Answer : 21,730.59.

Calculation:

Information provided in the question:

Original Cost 200000
Book Value 137000
Annual Appreciation Rate 6%
Sales Commission 5%
Tax on recaptured depreciation 25%
Tax on any capital gains 15%
Outstanding Mortgage Balance 155,000
Year Value Calculation
1 212000.00 [=200000+(200000*6%)]
2 224720.00 [=212000+(212000*6%)]
3 238203.20 [=224720+(224720*6%)]
4 252495.39 [=238203.2+(238203.2*6%)]
Forecast Sale Price: 252495.39
Selling Commission: 12624.7696 [=252495.39*5%]

Tax Payable on the Proposed Sale:

Forecast Sale Price 252495.39
Less: Selling Commission 12624.76
Less: Book Value 137000
Total Gain on Sales 102870.62
Capital Gain 39870.62 [=252495.39-12624.76-200000]
Recaptured Depreciation 63000 [=200000-137000]
Tax on Recaptured Depreciation 15750 [=63000*25%]
Tax on Capital Gain 5980.59 [=39870.62*15%]
Taxes Payable 21730.59

b.

Answer: 63,140.03

Calculation:

After-Tax Net Proceeds (CFR4)= 63140.03 [=252495.39-12624.76-155000-21730.59]